# CogCanvas: Compression-Resistant Cognitive Artifacts for Long LLM Conversations

### v1

@@ Abstract
On LoCoMo benchmark, CogCanvas achieves 34.7% overall accuracy, outperforming RAG (25.6%, +9.1pp) and GraphRAG (13.7%, +21.0pp). The advantage is most pronounced on temporal reasoning: 31.5% vs. 9.3% (RAG) and 5.0% (GraphRAG)—a +530% relative improvement. On multi-hop causal reasoning, CogCanvas achieves 81.0% pass rate vs. 40.0% for GraphRAG (+41.0pp).
@@ 1  Introduction
  • We provide a training-free, plug-and-play baseline that achieves strong performance on controlled benchmarks (97.5% recall, +78.5pp vs. summarization), with decisive advantages on complex reasoning: +41.0pp over GraphRAG on multi-hop reasoning (81.0% vs. 40.0%) and +530% relative improvement on temporal reasoning (31.5% vs. 5.0% GraphRAG), offering practitioners an immediately-deployable alternative without fine-tuning overhead.
Experimental evaluation demonstrates CogCanvas’s effectiveness across controlled benchmarks and real-world datasets. Our approach achieves 97.5% recall (+78.5pp vs. summarization) on information retention tasks, with two particularly striking results: (1) Multi-hop reasoning: 81.0% pass rate, outperforming GraphRAG by +41.0pp; (2) Temporal reasoning: 31.5% accuracy on LoCoMo temporal questions, representing a +530% relative improvement over GraphRAG (5.0%) and +239% over RAG (9.3%). These gains demonstrate that structured artifact preservation with temporal-aware construction provides decisive advantages for complex reasoning scenarios where standard baselines systematically fail.
@@ 2  Related Work
Our work intersects with several research directions: context management in LLMs, memory-augmented language models, graph-based knowledge organization, and conversation summarization.
2.1  Context Window Management
2.2  Memory-Augmented Language Models
MemGPT (now Letta) [packer2023memgpt] introduces a hierarchical memory system inspired by operating system design, requiring LLMs to learn explicit memory management operations. This architectural approach represents a fundamentally different paradigm from our inference-time retrieval, making direct comparison less meaningful.
2.3  Graph-Based Knowledge Organization
2.4  Conversation Summarization
2.5  Positioning of This Work
Table 1 summarizes the methodological distinctions between CogCanvas and related approaches. CogCanvas occupies a distinct point in the design space: lightweight enough for practical deployment, yet structured enough to resist the information loss inherent in alternative approaches.
@@ 3.3  Adaptive Injection
When generating a response to query $q$, CogCanvas retrieves and injects relevant artifacts. We employ a Hybrid Retrieval strategy that combines semantic similarity with lexical keyword matching. We compute the query embedding $e_{q}=\phi(q)$ and a keyword score for each canvas object. These are fused (e.g., via weighted sum or max-pooling for strong keyword matches) to generate a final relevance score. The top-$k$ objects based on this fused score form the candidate set $\mathcal{C}$.
@@
The complete injection pipeline operates transparently: from the LLM’s perspective, relevant context simply appears in the prompt, requiring no explicit memory management operations. Furthermore, the final stage of the pipeline incorporates Reasoning-aware Prompting, guiding the LLM to synthesize causal explanations from the retrieved artifacts, even if direct graph links are absent.
@@ 4.1  Experimental Setup
  • Summarization: Use GPT-4o-mini to compress conversation history into a summary, retaining recent 5 turns verbatim.
@@
  • GraphRAG: We include the official Microsoft GraphRAG implementation [edge2024graphrag] as a strong graph-based baseline. GraphRAG constructs a knowledge graph via entity extraction and community detection, then generates hierarchical community summaries for global queries. Our experiments use the default configuration with local search mode, which retrieves relevant entities and their community context.
@@ 4.4  Evaluation on Real-World Dataset (LoCoMo)
To valididate CogCanvas’s performance on real-world data, we evaluated it on the LoCoMo benchmark (ACL 2024), a dataset specifically designed for long-context memory and multi-hop reasoning. Unlike synthetic data, LoCoMo features natural conversational noise and complex information dependencies spanning over 300 turns.
@@
Native 7.9% 8.5% 2.8% 24.0%
Summ. 8.6% 10.3% 2.8% 22.9%
GraphRAG 13.7% 18.4% 5.0% 28.1%
RAG (k=10) 25.6% 39.7% 9.3% 38.5%
CogCanvas 34.7% 35.5% 31.5% 36.5%
$\Delta$ vs. best +9.1pp +22.2pp
Table 4: LoCoMo benchmark results (Categories 1–3, full evaluation).
@@
Table 4 summarizes the results. CogCanvas achieves 34.7% overall accuracy, outperforming all baselines including RAG (25.6%, +9.1pp) and GraphRAG (13.7%, +21.0pp).
The most striking result is on Temporal questions: CogCanvas achieves 31.5% compared to 9.3% (RAG) and 5.0% (GraphRAG), representing a +22.2pp absolute improvement (+239% relative vs. RAG, +530% vs. GraphRAG). This dramatic gap demonstrates that our temporal-aware graph construction and retrieval effectively captures time-sensitive information that other methods systematically lose.
On Single-hop retrieval, RAG leads (39.7%) due to its direct chunk matching, while CogCanvas (35.5%) achieves competitive performance while maintaining superior temporal reasoning. On Multi-hop questions, RAG achieves 38.5% through dense retrieval of related chunks, while CogCanvas (36.5%) demonstrates near-parity performance. This reveals an improved architectural balance: CogCanvas now achieves strong performance across all question types while specializing in temporal ordering and time-sensitive queries—the hardest aspect of long-context memory where all other methods fail catastrophically (2.8–9.3%).
Discussion on Absolute Performance. While CogCanvas demonstrates substantial advantages, the absolute scores on LoCoMo ($\sim$35%) reflect both the extreme difficulty of this benchmark and our more challenging evaluation setting: at the conversation midpoint, we trigger compression that discards early context entirely (retaining only recent 5 turns), then evaluate recovery from the compressed state.
CogCanvas outperforms all baselines: +21.0pp over GraphRAG (34.7% vs 13.7%) and +9.1pp over RAG (34.7% vs 25.6%). The advantage is most pronounced on Temporal questions (+22.2pp over RAG), validating CogCanvas as an effective solution for practitioners who need robust time-sensitive reasoning over compressed memory.
@@ 4.7  Discussion
The experimental results support our central hypothesis: structured artifact extraction and semantic retrieval provide compression-tolerant memory that outperforms lossy summarization. Three observations merit emphasis:
@@
Artifacts vs. Entities. The performance gap between CogCanvas and GraphRAG [edge2024graphrag] is instructive. GraphRAG achieves a respectable 43.2% pass rate through its sophisticated community detection and summarization pipeline. However, CogCanvas outperforms GraphRAG by +20.8pp. The key insight: while GraphRAG’s community summaries capture corpus-level themes effectively, they inevitably “compress away” the fine-grained causal reasoning (e.g., “…chose PostgreSQL because of JSONB support”) that our multi-hop questions target. CogCanvas succeeds precisely because it treats the verbatim artifact (Chunk + Type) as the atomic unit of memory, preserving the “connective tissue” of reasoning that summarization discards.
Resilience to Scale. CogCanvas effectively functions as a Session-Level Dynamic RAG system. By treating conversation history as a database of structured artifacts, it creates a persistent memory layer where retrieval accuracy is driven by query relevance rather than temporal proximity. For a retrieval system, finding a specific detail among 1,000 artifacts is computationally similar to finding it among 10; thus, it avoids the compression bottlenecks inherent in summarization.
@@
Exact preservation matters. The substantial exact match advantage (93.0% vs. 89.5% for RAG) demonstrates that CogCanvas systematically preserves specifics that other approaches lose. This gap illustrates the core distinction between distillation and dilution: summarization compounds errors through recursive decay—each compression pass further “waters down” the original semantics. CogCanvas instead performs semantic atomization, extracting crystallized knowledge units that remain intact regardless of subsequent context growth. For applications requiring faithful constraint adherence (coding assistants, technical support), this difference is practically significant.
Semantic retrieval is efficient. Top-5 retrieval achieves near-ceiling performance with modest token overhead ($\sim$250 tokens).
@@ 4.8  Error Analysis
The high proportion of temporal markers (14.9%) directly explains why temporal questions remain challenging (31.5% accuracy) despite dramatic improvements over baselines (5.0–9.3%). This suggests targeted improvements: explicit date/time extraction, temporal attribute modeling in artifacts, and retrieval strategies that prioritize temporal proximity for time-sensitive queries.
Performance by question type. Single-hop questions achieve 35.5% accuracy, multi-hop questions 36.5%, while temporal questions achieve 31.5%. Notably, our temporal-aware graph construction dramatically improves temporal reasoning compared to baselines (31.5% vs. 4–6% for all others), though the absolute performance remains limited due to extraction gaps in temporal markers.
@@ 4.9  Why Temporal Works
The most striking result from our LoCoMo evaluation is CogCanvas’s dramatic performance gain on temporal questions, achieving 31.5% accuracy compared to 4–6% for all baselines (+530% relative vs. GraphRAG, +239% vs. RAG). This section delves into the mechanisms underlying this significant improvement.
@@
    • CogCanvas’s Success: CogCanvas accurately retrieved the date, providing “May 7, 2023.” This was achieved by leveraging temporal-heuristic edges that linked the event to its precise timestamp in the original conversation, demonstrating its ability to preserve and retrieve granular temporal information.
@@
    • CogCanvas’s Success: CogCanvas precisely identified the date as “8 May 2022.” The temporal-heuristic edges enabled direct retrieval of the exact date, highlighting CogCanvas’s superior capability in retaining and reconstructing time-sensitive facts.
@@
To quantify the individual contributions of CogCanvas’s components, we conduct a systematic ablation study on the LoCoMo benchmark. We evaluate three configurations: (1) Baseline: graph construction with semantic retrieval only; (2) + CoT: adding chain-of-thought reasoning prompts; (3) Full System: adding temporal-heuristic edges and hybrid retrieval.
Configuration Overall 1-hop Temp. M-hop
Baseline (Graph + Semantic retrieval) 13.1% 12.8% 13.1% 11.5%
+ CoT prompting 29.9% 30.5% 27.1% 29.2%
Full System (+ Temporal + Hybrid) 34.7% 35.5% 31.5% 36.5%
Table 5: Ablation study on LoCoMo benchmark (Categories 1–3). Three configurations are evaluated: (1) Baseline uses graph construction with semantic retrieval; (2) + CoT adds chain-of-thought reasoning prompts (+16.8pp overall); (3) Full System additionally incorporates temporal-heuristic edges and hybrid retrieval (+4.8pp).
@@
  • CoT is the primary driver. The chain-of-thought prompting strategy contributes the majority of performance gain (+16.8pp, representing 77.8% of the total 21.6pp improvement). This demonstrates that guiding the LLM to synthesize causal explanations from discrete memory nodes is critical for complex reasoning tasks.
  • Full System provides additional gains. The combination of temporal-heuristic edges and hybrid retrieval adds a further +4.8pp overall. Notably, the Full System shows particularly strong gains on temporal questions (27.1%→31.5%, +4.4pp) and multi-hop reasoning (29.2%→36.5%, +7.3pp), suggesting these components provide targeted benefits for complex, time-sensitive queries.
  • Synergistic design matters. The performance improvement from Baseline (13.1%) to Full System (34.7%) represents a +21.6pp gain, confirming that no single component achieves strong performance in isolation. The full architecture—structured graph extraction, reasoning-aware prompting, temporal modeling, and hybrid retrieval—works synergistically to enable effective long-context memory.
@@
The consistently low performance of baselines (RAG, GraphRAG, Summarization) on temporal questions (4–6% accuracy) can be attributed to several factors:
@@
In contrast, CogCanvas’s verbatim-grounded artifact extraction combined with ‘temporal-heuristic edges‘ directly addresses these limitations by preserving explicit temporal markers and linking them structurally within the knowledge graph.

### v2

# CogCanvas: Verbatim-Grounded Artifact Extraction for Long LLM Conversations

@@ Abstract
On the LoCoMo benchmark (all 10 conversations from the ACL 2024 release), CogCanvas achieves the highest overall accuracy among training-free methods (32.4%), outperforming RAG (24.6%) by +7.8pp, with decisive advantages on complex reasoning tasks: +20.6pp on temporal reasoning (32.7% vs. 12.1% RAG) and +1.1pp on multi-hop questions (41.7% vs. 40.6% RAG). CogCanvas also leads on single-hop retrieval (26.6% vs. 24.6% RAG). Ablation studies reveal that BGE reranking contributes +7.7pp, making it the largest contributor to CogCanvas’s performance.
@@ 1  Introduction
  • We provide a training-free, plug-and-play baseline that achieves strong performance on controlled benchmarks (97.5% recall, +78.5pp vs. summarization) and the highest accuracy among training-free methods on LoCoMo (32.4% overall, +7.8pp vs. RAG), with decisive advantages on complex reasoning tasks: +20.6pp on temporal reasoning and +1.1pp on multi-hop questions. CogCanvas leads across all question types including single-hop retrieval (26.6% vs. 24.6%).
Experimental evaluation demonstrates CogCanvas’s effectiveness across controlled benchmarks and real-world datasets. Our approach achieves 97.5% recall (+78.5pp vs. summarization) on information retention tasks. On the LoCoMo benchmark (all 10 conversations from the ACL 2024 release), CogCanvas achieves the highest accuracy among training-free methods (32.4%), outperforming RAG (24.6%) by +7.8pp overall, with decisive advantages on complex reasoning—temporal questions (+20.6pp: 32.7% vs. 12.1%) and multi-hop questions (+1.1pp: 41.7% vs. 40.6%). CogCanvas leads across all question types including single-hop (26.6% vs. 24.6%).
@@ 2  Related Work
Our work intersects with context management, memory-augmented LLMs, graph-based retrieval, and conversation summarization. Context window management approaches include sparse attention [child2019sparsetransformer, beltagy2020longformer] and retrieval-augmented methods [borgeaud2022retro, izacard2022atlas]; we focus on what to preserve rather than extending capacity.
Memory-augmented LLMs range from hierarchical systems to heavily-optimized approaches. MemGPT [packer2023memgpt] pioneered OS-inspired virtual memory management for LLMs with self-editing memory tools; the original project has since evolved into Letta, [^1] https://letta.ai a production framework extending the core architecture with filesystem-based memory storage. Letta’s recent evaluation on LoCoMo reported 74.0% accuracy using GPT-4o mini [letta2024memory], demonstrating the effectiveness of stateful, database-backed memory systems—though this comes at the cost of requiring persistent infrastructure (PostgreSQL, Docker). EverMemOS [evermemos2025] achieves 92.3% on LoCoMo through categorical memory extraction and fine-tuning, while EMem [emem2024] reaches $\sim$78% via task-specific optimization. CogCanvas targets a complementary niche as a training-free baseline, prioritizing deployment simplicity over peak performance.
Graph-based methods like GraphRAG [edge2024graphrag] excel at static documents but struggle with dynamic conversational state. Summarization [xu2021dialogueSummary, chen2021dialogsum] operates on lossy compression, systematically losing specific details. See Appendix G for extended discussion.
@@ 3.2  Graph Construction
To capture implicit entities that may be missed in a single pass, we employ gleaning—a two-pass extraction strategy inspired by LightRAG [guo2024lightrag]. The second pass reviews the initial extraction results and specifically targets: (1) entities referred to by pronouns, (2) omitted subjects, (3) implicit causal relationships, and (4) temporal expressions. The results are merged and deduplicated, improving extraction recall by capturing nuanced information.
@@ 3.3  Adaptive Injection
When generating a response to query $q$, CogCanvas retrieves and injects relevant artifacts through a multi-stage pipeline.
Hybrid Retrieval with Adaptive Top-k.
We employ a Hybrid Retrieval strategy that combines semantic similarity with lexical keyword matching. We compute the query embedding $e_{q}=\phi(q)$ and a keyword score for each canvas object. These are fused to generate a final relevance score.
Critically, we use adaptive top-k selection based on question complexity: multi-hop questions (containing causal indicators like “because”, “after”, “why did”) retrieve $k=15$ objects; temporal questions (“when”, “how long”) retrieve $k=12$; simple factual questions use $k=10$. This allocates more context to complex queries while avoiding noise for simple ones.
Two-Stage Retrieval with Reranking.
To improve precision, we employ a two-stage retrieval process: (1) coarse retrieval fetches top-20 candidates using the hybrid scorer, then (2) a neural reranker (BGE-reranker) re-scores candidates based on query-document relevance. The top-$k$ objects after reranking form the final candidate set $\mathcal{C}$.
@@
The complete injection pipeline operates transparently: from the LLM’s perspective, relevant context simply appears in the prompt, requiring no explicit memory management operations. Furthermore, the final stage incorporates Reasoning-aware Prompting, guiding the LLM to synthesize causal explanations from the retrieved artifacts, even if direct graph links are absent.
@@ 4.1  Experimental Setup
  • Summarization: Use GPT-4o-mini to compress conversation history into a summary, retaining recent 5 turns verbatim. This baseline captures MemGPT’s [packer2023memgpt] core hierarchical memory mechanism.
@@
  • GraphRAG: We include the official Microsoft GraphRAG implementation [edge2024graphrag] as a graph-based baseline. GraphRAG constructs a knowledge graph via entity extraction and community detection, then generates hierarchical community summaries for global queries. Our experiments use the default configuration with local search mode (chunk size=800, overlap=100, entity types: person/technology/decision/fact/organization/concept, max_gleanings=1, community_level=2). We did not tune these parameters for conversational memory, which may underestimate GraphRAG’s potential on this task—see Limitations for discussion.
Note on Letta (MemGPT).
We attempted to include Letta (the production evolution of MemGPT [packer2023memgpt]) as a direct baseline. However, several technical barriers prevented controlled comparison:
  1. 1.
    Code availability: Letta’s LoCoMo evaluation code from the archived letta-leaderboard repository has been migrated to letta-evals, which does not currently support LoCoMo (GitHub issue #3115222https://github.com/letta-ai/letta/issues/3115).
  2. 2.
    Infrastructure requirements: Letta requires PostgreSQL database and Docker environment for persistent memory management, introducing non-deterministic factors (database state, API rate limiting) that complicate reproducible benchmarking.
  3. 3.
    Architectural divergence: Letta’s filesystem-based approach (storing conversations as files with semantic search) represents a fundamentally different paradigm from our stateless artifact extraction, making direct feature-level comparison less meaningful.
Nevertheless, we reference Letta’s official LoCoMo result (74.0% with GPT-4o mini [letta2024memory]) as an upper bound for stateful, database-backed systems in our discussion (§4.7.1).
@@ 4.4  Evaluation on Real-World Dataset (LoCoMo)
To validate CogCanvas’s performance on real-world data, we evaluated it on the LoCoMo benchmark (ACL 2024), a dataset specifically designed for long-context memory and multi-hop reasoning. Unlike synthetic data, LoCoMo features natural conversational noise and complex information dependencies spanning over 300 turns.
@@
Native 4.9% 5.3% 0.3% 18.8%
Summ. 5.6% 5.3% 0.6% 22.9%
GraphRAG 10.6% 12.8% 3.1% 29.2%
RAG (k=10) 24.6% 24.6% 12.1% 40.6%
CogCanvas 32.4% 26.6% 32.7% 41.7%
$\Delta$ vs. RAG +7.8pp +2.0pp +20.6pp +1.1pp
Table 4: LoCoMo benchmark results (Cat 1/2/3 only, all 10 conversations from the ACL 2024 release, binary LLM Judge). CogCanvas achieves the highest overall accuracy (32.4%) among training-free methods, leading across all question types with particularly strong temporal reasoning (+20.6pp vs. RAG).
@@
Table 4 summarizes the results using binary LLM-as-judge evaluation on official LoCoMo categories (single-hop, temporal, multi-hop). CogCanvas achieves the highest overall accuracy among training-free methods (32.4%), outperforming RAG (24.6%) by +7.8pp, while demonstrating advantages across all question types.
The most striking result is on Temporal questions: CogCanvas achieves 32.7% compared to 12.1% (RAG), representing a +20.6pp improvement. This substantial gap demonstrates that our verbatim artifact extraction effectively preserves time-sensitive information that chunk-based retrieval systematically loses. On Multi-hop questions, CogCanvas leads with 41.7% vs. 40.6% (RAG), a +1.1pp advantage, showing that graph-based expansion successfully connects related artifacts across conversation turns.
On Single-hop retrieval, CogCanvas also leads (26.6% vs. 24.6%), demonstrating that our enhanced retrieval parameters improve recall without sacrificing precision. The key insight is that CogCanvas provides consistent advantages across all question types, with particularly decisive gains on temporal reasoning (+20.6pp) where long-context memory matters most.
Discussion. This pattern validates our core hypothesis: structured artifact preservation provides decisive advantages for complex reasoning scenarios. CogCanvas’s overall lead (32.4% vs. RAG 24.6%) confirms that our approach excels across realistic long-context applications. Both retrieval-based methods substantially outperform GraphRAG (10.6%) and summarization (5.6%), confirming that retrieval-based approaches dominate lossy compression for long-context memory.
@@ 4.7  Discussion
The experimental results reveal a nuanced picture that refines our initial hypothesis. We organize our discussion around the fundamental trade-off between stateful and stateless memory architectures.
4.7.1  Stateful vs. Stateless Memory: A Trade-off Analysis
Our results reveal a fundamental trade-off between stateful and stateless memory architectures for long conversations:
Stateful Systems (e.g., Letta, EverMemOS).
These methods maintain persistent memory state across interactions, typically backed by databases or specialized storage. Letta achieves 74.0% on LoCoMo [letta2024memory], while EverMemOS reaches 92.3% [evermemos2025]. Their advantages include:
  • Superior recall: Persistent storage enables exhaustive memory retention without context length constraints.
  • Incremental learning: Agents can accumulate knowledge over extended sessions.
However, these benefits come at significant cost:
  • Infrastructure overhead: Requires databases (PostgreSQL), containerization (Docker), and server management.
  • Training requirements: Top performers like EverMemOS rely on learned compression models.
  • Deployment complexity: Non-trivial setup prevents rapid prototyping and integration into existing systems.
Stateless Systems (CogCanvas).
Our approach treats each query independently, extracting artifacts on-the-fly without persistent state. We achieve 32.4% on LoCoMo overall, outperforming all training-free baselines:
  • +7.8pp overall vs. RAG: Best training-free performance (32.4% vs. 24.6%).
  • +20.6pp on temporal queries: Graph-based retrieval with time-awareness outperforms RAG (32.7% vs. 12.1%).
  • +1.1pp on multi-hop reasoning: Verbatim grounding preserves logical chains that summarization destroys (41.7% vs. 40.6%).
  • Zero infrastructure: Drop-in integration via standard LLM APIs.
  • Training-free: No model fine-tuning or data collection.
When to Use Each Approach.
Our findings suggest:
  • Letta/stateful systems: Production deployments with long-lived agents, where infrastructure investment is justified by sustained interactions.
  • CogCanvas/stateless systems: Rapid prototyping, research experiments, or applications requiring portability and simplicity, especially when queries involve temporal reasoning or causal chains.
The 41.6pp gap between CogCanvas (32.4%) and Letta (74.0%) quantifies the “cost of simplicity”—a trade-off that may be acceptable in resource-constrained or exploratory settings where immediate deployability without infrastructure investment is prioritized.
4.7.2  Task Complexity Determines Optimal Approach
CogCanvas achieves the highest overall accuracy (32.4%), demonstrating that structured artifact preservation provides general benefits across all question types. CogCanvas leads on single-hop retrieval (26.6% vs. 24.6%), temporal questions (32.7% vs. 12.1%, +20.6pp), and multi-hop questions (41.7% vs. 40.6%, +1.1pp). This suggests that enhanced retrieval with structured memory provides consistent advantages—connecting constraints to decisions, tracking temporal sequences, and synthesizing information across conversation segments.
@@ 4.7.2  Task Complexity Determines Optimal Approach
Artifacts vs. Entities. The performance gap between CogCanvas and GraphRAG [edge2024graphrag] remains instructive. GraphRAG (10.6%) struggles on conversational memory despite its sophisticated community detection, because its summarization pipeline “compresses away” the fine-grained details that memory tasks require. CogCanvas succeeds precisely because it treats the verbatim artifact as the atomic unit of memory, preserving the “connective tissue” of reasoning that summarization discards.
Practical Trade-offs. CogCanvas achieves overall leadership (32.4%) with consistent advantages across all question types. For applications requiring temporal reasoning (“when did X happen?”) or multi-hop inference (“why was X decided?”)—common in project planning, technical support, and collaborative work—CogCanvas offers substantial benefits. The +20.6pp temporal advantage is particularly significant, as temporal reasoning represents the hardest aspect of long-context memory where all non-structured methods fail catastrophically (0.3%–12.1%).
Toward Hybrid Strategies. Our results suggest a practical deployment pattern: query-complexity routing. Simple queries containing direct entity references (“What is X’s email?”) can be routed to efficient RAG retrieval, while queries with temporal markers (“when,” “before,” “after”), causal indicators (“why,” “because,” “led to”), or multi-entity references can be routed to CogCanvas. A lightweight classifier—even regex-based heuristics—could achieve this routing with minimal overhead. This hybrid approach would combine RAG’s efficiency on simple lookups with CogCanvas’s strength on complex reasoning, potentially achieving higher overall accuracy than either method alone.
Target Applications. CogCanvas is particularly suited for domains where temporal and causal reasoning are critical: (1) project planning conversations tracking decisions, deadlines, and dependencies; (2) technical support sessions where constraints accumulate across turns; (3) collaborative coding where preferences and requirements must be faithfully preserved; (4) multi-stakeholder discussions (as demonstrated in our GitHub RFC case study) where cross-party decisions need structured tracking.
@@
Exact preservation matters. The substantial exact match advantage on controlled benchmarks (93.0% vs. 89.5% for RAG) demonstrates that CogCanvas systematically preserves specifics that other approaches lose. For applications requiring faithful constraint adherence (coding assistants, technical support), this difference is practically significant.
@@ 4.8  Error Analysis
What Gets Missed? Qualitative analysis of extraction failures reveals systematic patterns in what the LLM extractor overlooks:
  • Implicit personal attributes (e.g., “Caroline is single,” relationship status inferred from context rather than stated directly)—the extractor prioritizes explicit statements over implicit facts.
  • Emotional preferences (e.g., “she felt anxious about…”)—affective states are often not recognized as KeyFacts worth preserving.
  • Casual temporal references (“last Tuesday,” “a few months ago”)—relative time expressions are harder to anchor than absolute dates.
  • Negations and constraints (“does NOT want X”)—negative preferences are sometimes missed or inverted during extraction.
These patterns suggest that extraction prompts could be improved by explicitly instructing the LLM to attend to implicit attributes, emotional states, and negative constraints.
Gleaning contribution. Our two-pass gleaning strategy (Section 3.2) partially addresses these gaps by explicitly targeting omitted subjects and implicit relationships in the second pass. However, our ablation study reveals that gleaning provides minimal benefit on LoCoMo, likely because LoCoMo’s explicit conversational style reduces the need for implicit entity resolution. Gleaning may be more beneficial for informal dialogue with heavy pronoun usage.
The high proportion of temporal markers (14.9%) directly explains why temporal questions remain challenging (32.7% accuracy) despite substantial improvements over baselines (0.3%–12.1%). This suggests targeted improvements: explicit date/time extraction, temporal attribute modeling in artifacts, and retrieval strategies that prioritize temporal proximity for time-sensitive queries.
Performance by question type. Single-hop questions achieve 26.6% accuracy, multi-hop questions 41.7%, while temporal questions achieve 32.7%. Notably, our verbatim-grounded extraction substantially improves temporal reasoning compared to RAG (32.7% vs. 12.1%, +20.6pp), though the absolute performance remains limited due to extraction gaps in temporal markers.
@@ 4.9  Why Temporal Works
The most striking result from our LoCoMo evaluation is CogCanvas’s substantial performance gain on temporal questions, achieving 32.7% accuracy compared to 12.1% for RAG and 3.1% for GraphRAG—a +20.6pp improvement over the strongest baseline. This section delves into the mechanisms underlying this significant improvement.
@@
    • CogCanvas’s Success: CogCanvas accurately retrieved the date, providing “May 7, 2023.” This was achieved through verbatim artifact extraction that preserved the exact timestamp as a KeyFact, enabling direct retrieval of granular temporal information that summarization-based methods discard.
@@
    • CogCanvas’s Success: CogCanvas precisely identified the date as “8 May 2022.” The verbatim-grounded extraction preserved the exact date as a structured artifact, enabling direct retrieval—highlighting CogCanvas’s superior capability in retaining time-sensitive facts that compression-based methods lose.
@@
To quantify the individual contributions of CogCanvas’s components, we conduct a systematic ablation study by removing each component from the full system and measuring the performance drop.
Configuration All 1-hop Temp. M-hop
Full System 32.4% 26.6% 32.7% 41.7%
– Gleaning 30.7% 26.6% 28.7% 45.8%
– Graph Expansion 25.8% 19.5% 29.0% 33.3%
– Reranking 20.9% 16.7% 23.1% 26.0%
Minimal (Graph only) 14.3% 11.7% 13.7% 24.0%
Table 5: Ablation study on LoCoMo (Cat 1/2/3, all 10 conversations, LLM Judge). Full system achieves 32.4%. Reranking is the largest contributor (–11.5pp when removed), followed by graph expansion (–6.6pp).
@@
  • Reranking is the largest contributor. Removing BGE reranking causes a –11.5pp drop (32.4% $\to$ 20.9%), making it the most impactful component. Reranking improves precision by scoring retrieval candidates with semantic relevance, particularly benefiting multi-hop questions (41.7% $\to$ 26.0%, –15.7pp).
  • Graph expansion provides substantial gains. Removing graph expansion reduces overall accuracy by –6.6pp (32.4% $\to$ 25.8%), with the largest impact on multi-hop questions (–8.4pp). This confirms that structural relationships between artifacts enable discovering connected information across turns.
  • Gleaning shows minimal impact on LoCoMo. Removing the two-pass extraction strategy slightly improves overall accuracy (+1.7pp: 30.7% vs. 32.4%), suggesting that LoCoMo’s explicit conversational style reduces the need for implicit entity resolution. Gleaning may be more beneficial for informal dialogue with heavy pronoun usage.
  • Combined system effects. The full system (32.4%) improves +18.1pp over the minimal baseline (14.3%), demonstrating strong synergy between components.
  • Temporal advantage stems from extraction. CogCanvas’s temporal advantage over baselines (32.7% vs. RAG’s 12.1%, +20.6pp) stems primarily from verbatim artifact extraction preserving explicit timestamps, rather than from temporal heuristic edges. The extraction mechanism ensures time-sensitive information is retained as grounded artifacts.
Retrieval Quality Analysis.
To verify that graph expansion improves retrieval quality rather than merely leveraging LLM reasoning capabilities, we conduct a retrieval-only evaluation that measures whether the retrieved context contains ground-truth keywords—without invoking the LLM for answer generation.
Config. Overall 1-hop Temp. M-hop
Full System 71.8% 87.3% 63.2% 55.0%
– Graph Exp. 25.8% 32.3% 20.5% 24.7%
Improvement +46.0pp +55.0pp +42.7pp +30.3pp
Table 6: Retrieval recall: percentage of ground-truth keywords found in retrieved context (no LLM answering). Graph expansion improves retrieval quality by +46pp.
As shown in Table 6, graph expansion increases retrieval recall from 25.8% to 71.8% (+46.0pp). This confirms that the performance gains in Table 5 stem from improved context retrieval, not from LLM answering ability. The graph structure enables retrieving related artifacts that would otherwise be missed by pure semantic similarity matching.
@@
The consistently low performance of baselines on temporal questions (RAG: 12.1%, GraphRAG: 3.1%, Summarization: 0.6%) can be attributed to several factors:
@@
In contrast, CogCanvas’s verbatim-grounded artifact extraction directly addresses these limitations by preserving explicit temporal markers as structured KeyFacts. Our ablation study (Table 5) confirms that the temporal advantage (32.7% vs. RAG 12.1%, +20.6pp) stems primarily from this extraction mechanism, which ensures time-sensitive information is retained as grounded artifacts. The key insight is that preserving temporal information during extraction matters more than connecting it through graph structure.